# Evidence of Scaling Regimes in the Hopfield Dynamics of Whole Brain Model

### v2

@@ Introduction
Several comments are in order. First, we observe a clear increasing trend of $\alpha$ with $\delta$, meaning that small decay lengths, i.e. localised connectivity, promotes irregular patterns. For instance, below $\delta\sim 4$ mm, the scaling exponent is basically zero, corresponding to complete randomness. This was checked independently, by running simulations in which the couplings $J_{ij}$ were randomized though keeping the same distribution of the elements of the matrix $J$ (flat orange line at the bottom). We were able to simultaneously randomize the couplings $J_{ij}$, and keep the same value distribution by randomly shuffling the $J_{ij}$ values. At the opposite end, with $\delta$=10 mm, we measure $\alpha$$\sim$1.2, which is still well below the smoothness threshold $\alpha_{S}$=2, but above the random walk value $\alpha_{R}$=1, and much above the Kolmogorov turbulence value $\alpha_{T}$=2/3, which is in turn larger than $\alpha_{D}$$\sim$1/2, the value obtained by Deco et al.  (?) at $\delta$$\sim$5.55 mm. Importantly, in the intermediate regime, around the physiological value $\delta$=5.55 mm, the present Hopfield model delivers $\alpha_{H}$$\sim$2/5, different but still close to Deco et al. value. A visual extrapolation of Fig. 2A indicates that the scaling regime would remain non-smooth ($\alpha<2$) even in the limit of an ”infinite-brain”, $\delta\to\infty$), the global brain ($d\sim 25\delta$) being pretty close to the infinite-brain limit. Note however that for $\delta>10$ mm we start getting cases in which $B(d)\sim B(0)$, so that the numerical calculation of the exponent diverges and becomes less and less accurate (a fully correlated signal with $B(d)=B(0)$ yields formally $\alpha\to\infty$.)
@@
Figure 3 shows the pair correlation for different values of $\delta$. For $\delta<5$ mm, the pair correlation shows a basically uncorrelated activity, except at short range. For $\delta=5.26$ mm, long-range pair correlations start to emerge and finally, for larger $\delta$, the long-range pair correlations become even stronger. In the limit of very large $\delta$, stationary states emerge in which the system freezes into a single fully-ordered state, with all nodes states either $+1$ or $-1$. It is important to consider that in the fully-ordered state with no noise $\alpha$ is infinite.
@@
From this figure, a smooth trend at increasing $\rho$ is observed, with the scaling exponent starting to display significant decrease above $\rho=\left[0.95,0.98\right]$, when just 2 or 5 percent of the nodes are left, corresponding to $d_{ij}>4.6\delta$. With $J_{th}=0.10$, corresponding to $d_{ij}>-\delta ln(0.1)\sim 0.46\delta$, the scaling exponent is basically halved, indicating a major loss of correlation. This is expected since $J_{th}=0.10$ cuts out all but shortest range interactions. The conclusion of this dilution analysis is that the brain response is carried almost entirely by interactions up to about five decay lengths, namely about $1/6$ of the size of the global brain. This sounds reasonable, as it strikes a plausible compromise between short-range ($d<4\delta$) and long-range ($d>6\delta$) interactions.
It is interesting to observe that a “solid” (strongly correlated) brain, such as the one that results for $\delta>10$ mm, and low dilution $\delta$, would be too homogeneous to be able to store and process the amount of information required to function properly. On the opposite side, for too short connectivity length, the brain would behave like a “gas” of disconnected nodes, hence incapable of collective behavior, which is key to its proper functioning. It appears like the physiological length is achieving an optimal compromise between these two opposite, order-disorder, trends.
With a daring but conducive metaphor, we could speculate that the brain works in a sort of turbulent liquid-like state, although more spiky than actual turbulent fluids.
Changing The Number of Parcels
To assess the impact of the parcellation size on the scaling picture discussed in this paper, we gathered Schaefer 2018 Parcellations with different sizes $N$ and we run numerical simulations with $N_{r}=1000$ replicas and different values of $\delta$. This allowed us to inspect the relation between $\alpha$ and $\delta$ for different values of the size $N$. The result is the relation reported in Fig. 5, which shows a progressive regression of the transition regions towards smaller values of $\delta$, combined with a corresponding steepening of the transition.
@@
The sigmoid function captures the functional relation of the points reasonably well even though, as we discussed earlier on, the values $\alpha(\delta)$ do not fully converge in the top branch because for $\delta>10$ mm fully aligned configurations prevail, with correspondingly divergent values of $\alpha$.
Thus, we fitted the function in two steps: first, we estimated $\alpha_{\infty}$ on the plateau points for $N=800,900,1000$. Then, we obtained $\delta_{0}$ and $k$ by fitting the sigmoid function $\alpha(\delta)$ on all points using a nonlinear least squares fit. We finally obtained $\delta_{0}$ and $k$ as a function of $N$ via a linear least square fit in loglog scale, to deliver $\delta_{0}\sim N^{-0.379}$ and $k\sim N^{0.328}$, with a $R^{2}$-value $0.999$ and $0.925$, respectively
@@
Even if we cannot pin down such an optimal value, we expect it to lie between $N=10^{3}$, the largest parcellation presently available to us and $N<10^{6}$, which provides a plausible upper bound to the number of functioning ”modules” in the human brain (number of neurons, $10^{11}$, divided by the typical module size, $10^{5}$ neurons).
@@
Future experimental developments will allow us to explore finer parcellations with larger $N$ values and test the predictions made in this work. Furthermore, from these and other independent considerations  (?), it is plausible to speculate that for $N>10^{3}$, the connectivity may no longer be approximated by a decaying exponential.
@@
Conclusions
In conclusion, we have shown that an Hopfield recurrent neural network, informed by experimentally derived brain topology, recovers the scaling picture recently proposed by Deco  (?), according to which the process of information exchange within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows. Our analysis confirms the initial finding by  (?) in near-quantitative form, predicting a very similar scaling exponent ($2/5$ instead of $1/2$) for the same value of the connectivity decay length. Given the very different nature of the two mathematical models, this provides a significant hint at the robustness of the scaling picture.
@@ Conclusions
The Hopfield network model and the webXR parcels activity 3d representation will be available to access in open cloud servers ( e.g. H2IOSC data centers).
Acknowledgments
One of the authors (SS) wishes to acknowledge financial support from the Bertarelli Foundation during his stay at the Neurobiology Department of the Harvard Medical School. The authors wish to acknowledge illuminating discussions with John A. Assad, Haim Sompolinsky, and Bernardo Sabattini.
GG wishes to acknowledge financial support H2IOSC Project - Humanities and cultural Heritage Italian Open Science Cloud funded by the European Union – NextGenerationEU – NRRP M4C2 - Project code IR0000029 - CUP B63C22000730005.
Supplementary materials
Supplementary Text composed of the three paragraphs: i) the turbulent-like whole-brain model, ii) Amplitude of turbulence and scaling regime, iii) Structural connectivity.
Figure S1
Fig. 1. Structure factor $S_{2}(d)$ for $N=1000$ nodes and $N_{r}=1000$ realizations respectively at $\delta=5.55$ mm (A), and $\delta=5.88$ mm (B). The values $S_{2}(d)$ were binned over equal sized intervals. The dashed lines indicate different $\alpha$ values: $\alpha=2/3$ which corresponds to turbulence, $\alpha=1/2$, the value obtained in  (?), and $\alpha=2/5$. We estimate the scaling exponent $\alpha$ with a linear regression of $\log(S(d))\sim\alpha\log(d)+\beta$ in the interval range $[2.7-33.1]$ mm, as used in  (?), and we obtain $\alpha\approx 2/5$ for $\delta=5.55$ mm, and $\alpha\approx 2/3$ for $\delta=5.88$ mm.
Fig. 2. A) The blue line shows the average slope of $\alpha=\log S(d)/\log d$ for different $\delta$s, given the connectivity matrix $J_{ij}=\exp{(-d_{i,j}/\delta)}$. The horizontal dotted lines correspond to $\alpha_{RW}=1$ and $\alpha_{T}=2/3$, corresponding to random walk and homogeneous incompressible turbulence, respectively, while the dot-dashed horizontal line corresponds to Deco et al. “brain” value $\alpha_{D}=1/2$. The orange line shows the function $\alpha(\delta)$ with a connectivity matrix obtained by randomly shuffling all the weights $J_{ij}$, which results into a basically uncorrelated signal, indicating the importance of the metric structure of the weights. The plot stops at $\delta=10$ mm, a region where the signal is highly correlated ($\alpha\sim 1.2$), but still below the smooth regime marked by $\alpha=2$. Conversely, a brain below $4$ mm, would exhibit “chaotic” behaviour, with scaling exponents close to zero, i.e. no scaling regime. The bottom panels show the neuron activity mapped on the Schaefer’s cerebral cortical parcellation atlas coordinates for (B) $\delta=4$ (“chaotic”), (C) $\delta=5.99$, and (D) $\delta=6.66$ (random walk), respectively.
Fig. 3. Pair-correlation functions $B(d)$ for different $\delta$s. At small $\delta$ the signal is largely uncorrelated, while at increasing $\delta$, correlations start to emerge, although affected by significant amount of statistical fluctuations.
Fig. 4. The average slope of $\log S(d)$ for different dilution levels $\alpha(\rho)$, normalized to its threshold-free value $\alpha(0)$ at that given $\delta$. The dilution is measured as the fraction of disconnected pairs $i,j$, $J_{ij}$ elements with value 0. The second $y$-axis on top shows the cutoff distance in mm for which at a certain dilution $\rho$ the pairs at a larger or equal distance are disconnected. The differently colored curves indicate simulations with different $\delta$ values. The graph shows that within the measurement error, the system is not affected by the dilution of the edges, up to the removal of more than the $95\%$ of the connections, in descending order of distance. This shows that, in view of the exponential connectivity $J_{ij}=e^{-d_{ij}/\delta}$, the onset of collective patterns is sustained mostly by the close connections, within about $4\delta$.
Fig. 5. The average slope of $\alpha$ as a function of $\delta$ for Schaefer 2018 Parcellations with different sizes $N$. Given each size $N$, we fitted the points with the sigmoid functions in Eq. (7).

### v3

@@ 3 The Structure Factor $S_{2}(d)$
Several comments are in order. First, we observe a clear increasing trend of $\alpha$ with $\delta$, meaning that small decay lengths, i.e. localised connectivity, promotes irregular patterns. For instance, below $\delta\sim 4$ mm, the scaling exponent is basically zero, corresponding to complete randomness. This was checked independently, by running simulations in which the couplings $J_{ij}$ were randomized though keeping the same distribution of the elements of the matrix $J$ (flat orange line at the bottom). We were able to simultaneously randomize the couplings $J_{ij}$, and keep the same value distribution by randomly shuffling the $J_{ij}$ values. At the opposite end, with $\delta$=10 mm, we measure $\alpha$$\sim$1.2, which is still well below the smoothness threshold $\alpha_{S}$=2, but above the random walk value $\alpha_{R}$=1, and much above the Kolmogorov turbulence value $\alpha_{T}$=2/3, which is in turn larger than $\alpha_{D}$$\sim$1/2, the value obtained by Deco and Kringelbach (2020) at $\delta$$\sim$5.55 mm. Importantly, in the intermediate regime, around the physiological value $\delta$=5.55 mm, the present Hopfield model delivers $\alpha_{H}$$\sim$2/5, different but still close to Deco et al. value. A visual extrapolation of Fig. 2A seems to indicate that the scaling regime would remain non-smooth ($\alpha<2$) even in the limit of an "infinite-brain" ($\delta\to\infty$) the global brain ($d\sim 25\delta$) being pretty close to the infinite-brain limit. Note however that for $\delta>10$ mm we start getting cases in which $B(d)\sim B(0)$, so that the numerical calculation of the exponent diverges and becomes less and less accurate (a fully correlated signal with $B(d)=B(0)$ yields formally $\alpha\to\infty$.)
@@
Figure 3 shows the pair correlation for different values of $\delta$. For $\delta<5$ mm, the pair correlation shows a basically uncorrelated activity, except at short range. For $\delta=5.26$ mm, long-range pair correlations start to emerge and finally, for larger $\delta$, the long-range pair correlations become even stronger. In the limit of very large $\delta$, stationary states emerge in which the system freezes into a single fully-ordered state, with all nodes states either $+1$ or $-1$. As mentioned above in the fully-ordered state with no noise $\alpha$ is infinite.
Figure 3: Pair-correlation functions $B(d)$ for different $\delta$s. At small $\delta$ the signal is largely uncorrelated, while at increasing $\delta$, correlations start to emerge, although affected by significant amount of statistical fluctuations.
@@ 4 Thresholding the connections
Figure 4: The average slope of $\log S(d)$ for different dilution levels $\alpha(\rho)$, normalized to its threshold-free value $\alpha(0)$ at that given $\delta$. The dilution is measured as the fraction of disconnected pairs $i,j$, $J_{ij}$ elements with value 0. The second $y$-axis on top shows the cutoff distance in mm for which at a certain dilution $\rho$ the pairs at a larger or equal distance are disconnected. The differently colored curves indicate simulations with different $\delta$ values. The graph shows that within the measurement error, the system is not affected by the dilution of the edges, up to the removal of more than the $95\%$ of the connections, in descending order of distance. This shows that, in view of the exponential connectivity $J_{ij}=e^{-d_{ij}/\delta}$, the onset of collective patterns is sustained mostly by the close connections, within about $4\delta$.
From this figure, a smooth trend at increasing $\rho$ is observed, with the scaling exponent starting to display a significant decrease around $\rho=\left[0.95,0.98\right]$, when just 2 or 5 percent of the nodes are left, corresponding to $d_{ij}>4.6\delta$. With $J_{th}=0.10$, corresponding to $d_{ij}>-\delta ln(0.1)\sim 0.46\delta$, the scaling exponent is basically halved, indicating a major loss of correlation. This is expected since $J_{th}=0.10$ cuts out all but shortest range interactions. The conclusion of this dilution analysis is that the brain response is carried almost entirely by interactions up to about five decay lengths, namely about $1/6$ of the size of the global brain. This sounds reasonable, as it strikes a plausible compromise between short-range ($d<4\delta$) and long-range ($d>6\delta$) interactions.
It is interesting to observe that a “solid” (strongly correlated) brain, such as the one that results for $\delta>10$ mm, and low dilution $\delta$, would be too homogeneous to be able to store and process the amount of information required to function properly. On the opposite side, for too short connectivity length, the brain would behave like a “gas” of disconnected nodes, hence incapable of collective behavior, which is key to its proper functioning. It appears like the physiological length is achieving an optimal compromise between these two opposite, order-disorder, trends. With a daring but conducive metaphor, we could speculate that the brain works in a sort of turbulent liquid-like state, although more spiky than actual turbulent fluids.
5 Changing The Number of Parcels
To assess the impact of the parcellation size on the scaling picture discussed in this paper, we gathered Schaefer 2018 Parcellations (Schaefer et al., 2018) with different sizes $N$ and we run numerical simulations with $N_{r}=1000$ replicas and different values of $\delta$. This allowed us to inspect the relation between $\alpha$ and $\delta$ for different values of the size $N$. The result is the relation reported in Fig. 5, which shows a progressive regression of the transition regions towards smaller values of $\delta$, combined with a corresponding steepening of the transition.
Figure 5: The average slope of $\alpha$ as a function of $\delta$ for Schaefer 2018 Parcellations with different sizes $N$ (Schaefer et al., 2018). Given each size $N$, we fitted the points with the sigmoid functions in Eq. (7).
@@ 5 Changing The Number of Parcels
The sigmoid function captures the functional relation of the points reasonably well even though, as we discussed earlier on, the values $\alpha(\delta)$ do not fully converge in the top branch because for $\delta>10$ mm fully aligned configurations prevail, with correspondingly divergent values of $\alpha$. Thus, we fitted the function in three steps: first, we estimated $\alpha_{\infty}$ on the plateau points for $N=800,900,1000$. Then, we obtained $\delta_{0}$ and $k$ by fitting the sigmoid function $\alpha(\delta)$ on all points using a nonlinear least squares fit. We finally obtained $\delta_{0}$ and $k$ as a function of $N$ via a linear least square fit in loglog scale, and found $\delta_{0}\sim N^{-0.379}$ and $k\sim N^{0.328}$, with a $R^{2}$-value $0.999$ and $0.925$, respectively.
@@
Even if we cannot pin down such an optimal value, we expect it to lie between $N=10^{3}$, the largest parcellation presently available to us and $N<10^{6}$, which provides a plausible upper bound to the number of functioning “modules” in the human brain (number of neurons, $10^{11}$, divided by the typical module size, $10^{5}$ neurons).
@@
Future experimental developments will allow us to explore finer parcellations with larger $N$ values and test the predictions made in this work. Furthermore, from these and other independent considerations Kaiser (2014), it is plausible to speculate that for $N>10^{3}$, the connectivity may no longer be approximated by a decaying exponential.
@@
6 Conclusions
In conclusion, we have shown that an Hopfield recurrent neural network, informed by experimentally derived brain topology, recovers the scaling picture recently proposed by Deco and Kringelbach (2020), according to which the process of information exchange within the human brain shows spatially correlated patterns qualitatively similar to those displayed by turbulent flows. Our analysis confirms the initial finding by Deco and Kringelbach (2020) in near-quantitative form, predicting a very similar scaling exponent ($2/5$ instead of $1/2$) for the same value of the connectivity decay length. Given the very different nature of the two mathematical models, this provides a significant hint at the robustness of the scaling picture.
@@ 6 Conclusions
7 Acknowlegments
SS wishes to acknowledge financial support from the Bertarelli Foundation during his stay at the Neurobiology Department of the Harvard Medical School. GG wishes to acknowledge financial support H2IOSC Project - Humanities and cultural Heritage Italian Open Science Cloud funded by the European Union – NextGenerationEU – NRRP M4C2 - Project code IR0000029 - CUP B63C22000730005. The authors wish to acknowledge illuminating discussions with John A. Assad, Haim Sompolinsky, and Bernardo Sabattini.